### v1

# g-DPO: Scalable Preference Optimization for Protein Language Models

Abstract
Direct Preference Optimization (DPO) is an effective approach for aligning protein language models with experimental design goals. However, DPO faces a scalability bottleneck: the number of possible training pairs grows quadratically with the number of labeled sequences, leading to prohibitive training times even for modestly sized datasets. We introduce g-DPO, a framework that (i) uses sequence space clustering to prune redundant pairs while preserving training signal, and (ii) amortizes likelihood computations with group-based approximations. Across three protein engineering tasks, g-DPO maintains in-silico and in-vitro performance that is statistically indistinguishable from standard DPO, while converging $1.8$ to $3.7$ times faster, with greater gains expected as the size of the dataset increases.
1 Introduction
Protein language models (PLMs), trained with self-supervised learning on large sequence datasets, capture relevant structure and function signals [17, 33, 25, 14]. To use these models for protein engineering tasks, such as optimizing binding affinity, thermostability, or catalytic activity, they are typically fine-tuned on labeled experimental datasets, many of which naturally define preferences (e.g., higher fluorescence is better, lower toxicity is better).
Direct Preference Optimization (DPO) is a natural fit for this setting. DPO fine-tunes a language model using pairwise preferences, maximizing the likelihood of preferred sequences relative to dispreferred ones [24], without requiring a separate reward model as in RLHF [6]. Recent studies have applied DPO to protein sequence design [2, 32, 30, 37], however, unlike in NLP, where explicit human preferences are available [22, 3], experimental protein datasets often provide scalar labels. Constructing preferences by exhaustively comparing all samples is infeasible because the number of pairs grows quadratically with dataset size. Moreover, not all comparisons are equally informative. Distant comparisons in sequence space tend to collapse into coarse binary signals, while local comparisons capture subtle and often non-additive effects of a few mutations [12, 10], providing a more valuable learning signal, which is crucial in late-stage optimization.
We introduce group-DPO (g-DPO), a DPO framework for experimentally labeled protein data that addresses these challenges. This study contributes: (1) Scalable preference sampling: Sequence-space clustering prunes redundant comparisons and focuses training on informative local neighborhoods in input space. (2) Efficient training: Building on this clustering, we exploit union masking to amortize log-likelihood computations across groups of sequences, further improving convergence time. (3) Empirical validation: Across three protein mutational landscapes, g-DPO maintains the in-silico and in-vitro performance of standard DPO, while converging $1.8$ to $3.7$ times faster, with larger convergence time gains expected as the size of the dataset increases.
2 Related work: Constructing pairs for DPO with PLMs
Direct Preference Optimization (DPO). DPO fine-tunes language models directly from pairwise comparisons, avoiding the need for a separate reward model [24]. Given a distribution $\mathcal{D}$ of pairs of sequences $(y_{w},y_{l})$, where $y_{w}$ (winner) is preferred over $y_{l}$ (loser), a policy $\pi_{\theta}$ is optimized to increase the likelihood of sampling $y_{w}$ over $y_{l}$ under a regularization constraint formulated as a bound on the KL divergence between $\pi_{\theta}$ and a reference policy $\pi_{\mathrm{ref}}$. This objective is expressed as a loss function:
$\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta},\pi_{\mathrm{ref}})=-\mathbb{E}_{(y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\Big(\beta\log\frac{\pi_{\theta}(y_{w})}{\pi_{\mathrm{ref}}(y_{w})}-\beta\log\frac{\pi_{\theta}(y_{l})}{\pi_{\mathrm{ref}}(y_{l})}\Big)\right],$ (1)
where $\beta>0$ is the regularization hyperparameter and $\sigma$ is the sigmoid function. A complete derivation is provided in A.1. A primary challenge in applying DPO to PLMs is constructing a preference distribution $\mathcal{D}$ from scalar labels, as exhaustively pairing labeled sequences does not scale.
2.1 Preference sampling for DPO with PLMs
Output space partitioning.
A common strategy is to partition labeled sequences by thresholds defined by design objectives (e.g., "get sequences above $x$ stability") or quantiles, and assign sequences above as "preferred" and sequences below as "dispreferred" [20, 35, 34, 26, 36]. In terms of complexity, subsampling schemes, such as fixing $k$ losers per winner, can be used and scale effectively with dataset size. Although this aligns well for coarse optimization goals, treating all pairs within the same partition as equivalent reduces resolution among high-performing sequences, which is limiting in late-stage optimization.
Rank-space sampling.
Unlike thresholding, which collapses labels into binary partitions, rankings preserve relative relationships across all sequences. Methods such as RRHF [38] and PRO [28] show that training on rankings (or their induced pairs) leads to a more stable convergence in aligning LMs with human preferences. Applied to proteins, CtrlProt [15] shows that a rank-wise preference objective improves controllability over pairwise methods for multi-objective optimization. Widatalla et al. [32] evaluate a ‘gap level’ heuristic that down-samples pairs of sequences at fixed distances in rank space, and find that most gap levels have comparable performance to randomly sampling pairs. This suggests that not all pairs are equally informative and that randomly pruning pairs in the output space risks diluting the training signal.
Informative pairs.
Other methods have been proposed to prioritize preference pairs that provide a stronger training signal. Maru et al. [18] propose a hybrid sampling strategy that mixes global and local perturbations, selecting pairs with large predicted stability differences for the former and pairs ten mutations apart for the latter. They find no improvement over randomly selecting pairs. Beyond proteins, other strategies have been proposed to improve informativeness, such as stratified sampling to ensure coverage across score bins, oversampling rare but informative regions, or curriculum learning to gradually introduce harder comparisons as the model improves [13, 23, 7, 4, 11].
3 Methods: Group-DPO (g-DPO)
In NLP, preferences are inherently structured: multiple outputs are generated from the same prompt, and annotators decide which is better [22, 3]. This provides a shared context that makes likelihood ratios between outputs meaningful. On the other hand, comparing outputs from different prompts yields likelihoods ratios that are dominated by prompt differences rather than output preference. Proteins have a related dynamic: even when variants come from the same wild type, comparing distant variants conflates the effects of many mutations, collapsing into coarse signals, whereas local comparisons better capture subtle, non-additive effects [12, 10]. Our work, therefore, leverages proximity in sequence space to construct preference pairs, enabling the model to learn fine-grained mutational effects while reducing the number of training pairs.
Our g-DPO training framework is conducted in three stages. We begin with unsupervised fine-tuning of a PLM on evolutionarily related sequences of a wild type protein (evo-tuning) [1, 8]. Next, given an experimental mutant dataset of the relevant wild type, we apply union mask clustering, a greedy agglomerative procedure that groups sequences based on shared mutational positions.
Figure 1: Clustering of mutants with overlapping mutations.
Finally, we sample groups of sequences from each cluster and evaluate the DPO loss over all preferences within each group to further fine-tune our evo-tuned model. Although some assay variants could appear within the evo-tuning training data, this does not constitute data leakage as evo-tuning uses no assay labels. This unsupervised design choice is intended to provide general evolutionary context to the model, opposed to labeled samples for model alignment.
3.1 Union mask clustering
Given a set of aligned sequences $S=(s^{1},\ldots,s^{n}),\;s^{i}\in\mathcal{A}^{L}$, over the amino acid alphabet $\mathcal{A}$, define the union mask as $M(S)=\{i\in[\![L]\!]:\exists j,k\ \text{such that }s_{i}^{(j)}\neq s_{i}^{(k)}\}$, and let $m(S):=|M(S)|$ denote the size of the union mask. We cluster sequences with a greedy agglomerative procedure with a linkage that prefers merges that keep the union masks small:
  1. 1.
    Initialization. Let $\mathcal{C}\leftarrow\{\{s^{1}\},\ldots,\{s^{n}\}\}$. For each cluster $C\in\mathcal{C}$, store its union mask $M(C)$.
  2. 2.
    Linkage. For clusters $C_{i},C_{j}$, define the cost of merging $C_{j}$ into $C_{i}$ as
    $\phi(C_{i},C_{j})=m(C_{i}\cup C_{j})-m(C_{i})=\big|M(C_{i})\cup M(C_{j})\cup M(\{s,s^{\prime}\})\big|-m(C_{i}),$
    where $s\in C_{i}$ and $s^{\prime}\in C_{j}$ are arbitrarily chosen.
  3. 3.
    Greedy merge. Repeatedly select
    $(C_{p},C_{q})=\arg\min_{i\neq j}\phi(C_{i},C_{j}),$
    breaking ties by minimum $m(C_{i})$. Replace clusters $C_{p}$ and $C_{q}$ by $C_{p}\cup C_{q}$, and store the new union mask $M(C_{p}\cup C_{q})=M(C_{p})\cup M(C_{q})\cup M(\{s,s^{\prime}\})$.
  4. 4.
    Stopping. Stop when the next best merge would violate the union mask ratio threshold $\tau\in[0,1]$,
    $\min_{i\neq j}\phi(C_{i},C_{j})>\tau L.$
Time complexity.
Calculating $\phi(C_{i},C_{j})$ is $O(L)$, since $M(C_{i})$ and $M(C_{j})$ are stored, and only $M({s,s^{\prime}})$ must be computed. In each merging step, computing all costs $\phi$ is $O(n^{2}L)$. Using a heap over cluster pairs, the final time complexity is $O(n^{2}\log n+n^{2}L)$. When $n$ is large, we first coarse-cluster with MMseqs2 [29], which provides approximate linear-time clustering for large datasets, and then run union mask clustering independently within those buckets.
3.2 Group sampling
Log likelihood approximation.
[…]
Since union-mask clustering controls the union-mask size of clusters, we can extend this approximation beyond pairs. We uniformly sample without replacement groups of $g$ sequences from each cluster. We can then approximate the likelihood of each sequence in a group with a single forward pass. We then evaluate the DPO loss over all pairwise comparisons inside the group. Sampling continues until every sequence has appeared in at least one group, which defines one training epoch.
4 Experiments
(a) (a)
(b) (b)
(c) (c)
Figure 2: In-silico evaluation of g-DPO. (a) Spearman correlation ($\rho$) of evo-tuned, DPO, and g-DPO models across the three datasets. (b) Relative convergence speed of g-DPO compared to DPO, measured on a single NVIDIA A100 GPU, with $\Delta\rho$ indicating performance differences. (c) Predicted property distributions of sequences generated from the reference, DPO, and g-DPO models using beam search. Kolmogorov-Smirnov tests confirm that g-DPO and DPO yield nearly identical, and improved distributions over the reference.
Evaluating generative models for protein optimization is challenging and ultimately requires experimental evidence. Our goal is to test whether g-DPO (i) preserves or improves model quality relative to DPO while (ii) reducing training cost, and (iii) whether in-silico improvements translate to experimental outcomes. We evaluate g-DPO on experimentally measured variants derived from three wild type proteins:
  • Anti-SARS-CoV-2 VHH [9]: thermostability ($N=462$, mutation breadth 47.1%).
  • Trastuzumab scFv: expression ($N=76$, mutation breadth 13.1%).
  • Haloalkane dehalogenase (DhaA): thermostability ($N=474$, mutation breadth 40.3%).
We use ESM-2-650 [14] as our pre-trained model. Additional dataset and training details are provided in Appendix C.
4.1 In-silico validation
For each dataset, we report (i) rank correlation between PLL scores and experimental measurements on holdout test sets, which tests whether model likelihoods align with ground truth preferences; (ii) convergence time to early stopping, to measure training efficiency; and (iii) generative quality, measured by scoring sequences generated with beam search [31]. We score the generated sequences with independent predictors trained on the same experimental datasets.
Both DPO and g-DPO improve rank correlation relative to the evo-tuned reference model (Figure 2a) and shift the distribution of generated sequences toward better predicted function (Figure 2c).
Refer to caption Refer to caption
(a) (b)
Figure 3: In-vitro validation. Distribution of assay measurements for sequences designed with g-DPO and DPO on (a) thermostability of DhaA and (b) expression of Trastuzumab. Distributions show that both models yield comparable outcomes.
The distributions produced by DPO and g-DPO are statistically indistinguishable, indicating that g-DPO preserves generative quality while reducing training cost (Figure 2b). Statistical significance was assessed with two-sample KS tests, with full results reported in Table 2.
4.2 In-vitro validation
To test whether in-silico improvements translate to experimental outcomes, we validated a subset of generated sequences in the lab for two of the three optimization tasks: thermostability of DhaA and expression of Trastuzumab. Candidates were down-selected from the in-silico generation pool using a criterion that averages performance across the top-$k$ predicted sequences ($k=3$), following the Monte Carlo sampling strategy of DiscoBAX [16]. Assay results confirm that DPO and g-DPO yield comparable experimental outcomes, with no significant differences observed between their distributions.
4.3 Ablations
We ablate clustering ($\tau$) and grouping ($g$) to understand their individual effects and how they interact. All ablations are reported on the Anti-SARS-CoV-2 VHH dataset.
Figure 4: (a) Effect of clustering threshold $\tau$ on performance and training pairs. (b) Effect of grouping size $g$ on performance and convergence speed without clustering.
Clustering, no grouping.
We sweep the clustering threshold $\tau$ from $0.5\rightarrow 0.1$, where a smaller $\tau$ yields more clusters, reduces the number of training pairs and speeds up convergence. In Figure 4a with $g=2$, we see that as $\tau$ decreases to $\approx 0.3$, model performance remains unchanged. Beyond this threshold, performance declines as clusters become too tight and useful signal is lost. This indicates that many pairs are redundant and can be safely pruned in input space up to a moderate threshold ($\tau\approx 0.3$).
Grouping, no clustering.
Grouping amortizes likelihood computations by sharing a union mask and a single forward pass across $g$ sequences, but comes at a cost. When mutations span a large fraction of the sequence, union masks become too large, likelihoods are overly approximated, and model performance drops (Figure 4b). Hence, grouping alone is insufficient when the mutation span is high.
Clustering with grouping.
Combining clustering with grouping ($g=4$ in Figure 4a) yields the expected trade-off. With little to no clustering, grouping degrades model performance, as previously mentioned. As the clustering threshold increases, performance improves and eventually matches that of $g=2$ at $\tau\approx 0.3$, while convergence is significantly faster for $g=4$. Beyond $\tau\approx 0.3$, performance drops as clustering becomes too strict and valuable training signal is lost. This drop is less severe for $g=4$, probably because within-group preferences are processed in one batch, with coupled likelihoods and update steps, while with $g=2$, they can be split across batches. Practically, clustering controls the training signal (removing redundant pairs), and grouping controls efficiency; however, accurate grouping is dependent on a reasonable clustering level for the likelihood approximation to hold.
5 Conclusion, limitations and future work
We present g-DPO, a scalable variant of DPO for PLMs, that combines sequence-space clustering, to prune redundant preference pairs, with grouped likelihood amortization, to reduce computation. Across three protein engineering tasks, g-DPO matches DPO performance on both in-silico metrics and in-vitro results, while converging up to 3.7× faster. In practice, moderate clustering with grouping yields the best trade-off; excessive clustering removes training signal, and grouping alone can degrade performance when mutation span is large and likelihood approximations break down.
Our experiments span three mutational landscapes with 76-474 variants, representative of typical assay-driven protein engineering datasets. However, extending evaluation to larger and more diverse benchmarks (multiple wild types, broader mutation spans) would allow to further test the scalability and robustness of g-DPO. We expect the scalability advantage of g-DPO to become even more important in such settings. A natural extension is to move beyond a pairwise objective and explore rank-based objectives which could leverage the full ranking information of variants within clusters. Moreover, while this study focused on unimodal protein sequence PLMs, the clustering and grouping strategies are modality-agnostic. Extending g-DPO to multi-modal foundation models (e.g., sequence-structure or sequence-function) is a promising direction for future work.
Appendix A Preliminaries
A.1 Direct Preference Optimization (DPO) Overview
Given a pre-trained reference model $\pi_{\text{ref}}$, the goal of DPO is to learn a policy $\pi_{\theta}$ that maximizes the probability of generating preferred sequences over dispreferred ones.
Consider a dataset of $N$ pairwise rankings, $D=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, where $x^{(i)}$ is the masked input and $y_{w}^{(i)}$ and $y_{l}^{(i)}$ are the preferred and dispreferred outputs. The preference distribution $p^{*}$ can be expressed given a Bradley-Terry reward model $r^{*}$ and a parameterization of the score $s(x,y)=\exp(r^{*}(x,y))$,
$p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r^{*}(x,y_{w}))}{\exp(r^{*}(x,y_{w}))+\exp(r^{*}(x,y_{l}))}=\sigma(r^{*}(x,y_{w})-r^{*}(x,y_{l}))$
Under the RLHF framework, a parameterized reward model $r_{\Phi}(x,y)$ can be estimated to match the optimal reward function $\pi_{*}$, using the negative log-likelihood loss,
$L_{R}(r_{\Phi},D)=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\,[\log\sigma(r_{\Phi}(x,y_{w})-r_{\Phi}(x,y_{l}))]$
This reward model can then be used to optimize the optimal policy,
$\pi^{*}=\arg\max\mathbb{E}_{x\in D,y\sim\pi_{\theta}}[r_{\Phi}(x,y)]-\beta D_{\text{KL}}[\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x)]$
The KL term is there to ensure that the outputs from the learned policy do not diverge too much from the original policy, as we assume that the fine-tuned base model already produces reliable outputs. Using Gibbs’ inequality, we can derive the optimal solution to this optimization problem,
$\pi^{*}(y\mid x)=\frac{1}{Z(x)}\pi_{\text{ref}}(y\mid x)e^{\tfrac{1}{\beta}r_{\Phi}(x,y)}$
where $Z(x)=\sum_{y}\pi_{\text{ref}}(y\mid x)e^{\tfrac{1}{\beta}r_{\Phi}(x,y)}$, an untractable term. We can reorganize this term to solve for $r_{\Phi}$ or $r^{*}$ with its corresponding policy $\pi^{*}$, which we can plug back into (2) to obtain,
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(y_{w},y_{l})\sim D}\left[\log\sigma\big(\beta\log\tfrac{\pi_{\theta}(y_{w})}{\pi_{\text{ref}}(y_{w})}-\beta\log\tfrac{\pi_{\theta}(y_{l})}{\pi_{\text{ref}}(y_{l})}\big)\right]$
where $\sigma$ is the sigmoid function and $\beta$ is a scaling parameter.
This objective encourages the policy to increase the likelihood of preferred sequences relative to dispreferred ones, while the reference model serves as a regularizer to prevent the policy from deviating too far from the original distribution.
Appendix B g-DPO Details
B.1 Union mask clustering algorithm
Algorithm 1 Greedy Union-Mask Clustering
1:Sequences $\{s^{1},\dots,s^{n}\},s^{i}\in\mathcal{A}^{L}$; length $L$; union-mask ratio threshold $\tau\in[0,1]$.
2:A set of clusters $\mathcal{C}$.
3:Initialization: $\mathcal{C}\leftarrow\big\{\{s^{1}\},\dots,\{s^{n}\}\big\}$.
4:for all $C\in\mathcal{C}$ do
5:  Store its union mask $M(C)$ and its size $m(C)=|M(C)|$.
6:end for
7:function $\phi$($C_{i},C_{j}$) $\triangleright$ Merge cost of $C_{j}$ into $C_{i}$
8:  Choose arbitrary $s\in C_{i}$, $s^{\prime}\in C_{j}$.
9:  return $|\,M(C_{i})\cup M(C_{j})\cup M(\{s,s^{\prime}\})\,|-m(C_{i})$.
10:end function
11:while true do
12:  $(p,q)\leftarrow\arg\min_{i\neq j}\big(\phi(C_{i},C_{j})\big)$ $\triangleright$ Tie-break: smaller $m(C_{i})$
13:  $c^{\star}\leftarrow\phi(C_{p},C_{q})$
14:  if $c^{\star}>\tau L$ then $\triangleright$ Stopping rule
15:   break
16:  end if
17:  Choose arbitrary $s\in C_{p}$, $s^{\prime}\in C_{q}$.
18:  $C_{\text{new}}\leftarrow C_{p}\cup C_{q}$
19:  $M(C_{\text{new}})\leftarrow M(C_{p})\cup M(C_{q})\cup M(\{s,s^{\prime}\})$
20:  $m(C_{\text{new}})\leftarrow|M(C_{\text{new}})|$
21:  $\mathcal{C}\leftarrow\big(\mathcal{C}\setminus\{C_{p},C_{q}\}\big)\cup\{C_{\text{new}}\}$
22:end while
23:return $\mathcal{C}$
B.2 Single forward pass log-likelihood approximation algorithm
The DPO loss requires log-likelihood differences between sequences. By constructing a union mask for a group $G$, all sequences within $G$ are evaluated under the same masked context, making their likelihoods directly comparable. This ensures that the log-likelihood ratios used in the loss are valid. Importantly, the approximation only holds for sequences scored under the same union mask.
Algorithm 2 Union-mask $L$ approximation
1:Cluster $C\subset\mathcal{C}$; group $G\subset C$ of size $g$; MLM model $\pi$; mask token [MASK]
2:Per-sequence approximate likelihoods $\{\widetilde{\log p_{\pi}}(y)\}_{y\in G}$
3:function UnionMask($G$)
4:  return $D\leftarrow\{\,p\in[\![L]\!]\mid\exists\,y,y^{\prime}\in G:\;y_{p}\neq y^{\prime}_{p}\,\}$
5:end function
6:function Mask($y,D$)
7:  $x\leftarrow y$;  $x_{p}\leftarrow\texttt{[MASK]}$ for all $p\in D$
8:  return $x$
9:end function
10:function Likelihoods($G,\pi$)
11:  $D\leftarrow\textsc{UnionMask}(G)$
12:  Choose arbitrary $y\in G$
13:  $x\leftarrow\textsc{Mask}(y,D)$
14:  $\mathrm{logits}\leftarrow\pi(x)$ $\triangleright$ single forward pass; returns token logits for all positions
15:  for all $y\in G$ do
16:   $\widetilde{\log p_{\pi}}(y)\leftarrow\sum_{p\in D}\log\mathrm{softmax}(\mathrm{logits}_{p})[\,y_{p}\,]$
17:  end for
[…]
19:end function
B.3 Training speedup
[…]
Appendix C Experiment details
C.1 Training details
We start with pre-trained ESM-2-650 (weights are open-sourced) [14], which we fine-tune on evolutionarily related sequences of our wild type sequence for each corresponding dataset [1, 8]. We follow the evo-tuning framework defined by [1]. Then, this model is further fine-tuned using the g-DPO framework with SGD using a learning rate of $7\times 10^{-4}$ and no weight decay, with a 300-step linear warmup. The DPO loss is computed with $\beta=0.04$ with a batch size of 64. Sequences were clustered with a maximum union mask size of $0.3L$, and groups of $g=4$ were sampled from each cluster. The loss was applied over all within-group pairs. This was the final configuration, though we ran hyperparameter sweeps and expect the optimal group size to depend on dataset scale and clustering. Validation loss was monitored every 250 steps, with checkpointing on validation loss and relative early stopping on validation loss that checks there has not been more than a 1% improvement with a patience of 3 validation intervals. Training was performed on a single NVIDIA A100 GPU.
C.2 Datasets
(a) DhaA
(b) Anti-SARS-CoV-2 VHH
(c) Trastuzumab scFv
Figure 5: Dataset distributions.
anti-SARS-CoV-2 VHH: The wild type[9] already expresses well and has a binding affinity in the single-digit nanomolar range.
DhaA: Haloalkane dehalogenase catalyzes the hydrolysis of halogenated compounds by cleavage of the carbon-halogen bond.
Trastuzumab scFv: An scFv version of the cancer drug Trastuzumab, which expresses poorly and has a binding affinity of 3–5 nanomolar.
Protein Function NN Mutation breadth WT measurement
anti-SARS-CoV-2 VHH Thermostability 462 47.1% 58.67 C
Trastuzumab scFv Expression 76 13.1% 0.28
Haloalkane dehalogenase Thermostability 474 40.3% 45.09 C
Table 1: Optimized function, dataset size (NN), mutation breadth (% of positions mutated), and wild type measurement for datasets used in experiments.
Appendix D Additional results
D.1 In-silico results
We report Kendall’s $\tau$ as a complementary rank correlation metric to Spearman $\rho$. While Spearman captures monotonic correlation, Kendall’s $\tau$ measures the fraction of concordant versus discordant pairs and is therefore more sensitive to local ordering differences. Formally, it is defined as
$\tau=\frac{(\#\text{concordant pairs})-(\#\text{discordant pairs})}{\#\text{total pairs}}.$
We report additional results with Kendall’s $\tau$ in Figure 6 and Figure 7. The consistent trends across Spearman $\rho$ and Kendall’s $\tau$ confirm that our conclusions are robust to the choice of ranking measure.
Figure 6: Kendall tau ($\tau$) correlation between PLL scores and experimental measurements on holdout test sets for the evo-tune, DPO, and g-DPO models.
Figure 7: Effect of clustering threshold $\tau$ on model performance and number of effective training pairs. Moderate clustering prunes redundant pairs without loss of performance.
D.2 Statistical significance tests
We include two-sample Kolmogorov-Smirnov (KS) tests on the predicted property distributions of generated sequences (see Table 2). While the KS tests indicate statistically significant differences between DPO and g-DPO, the effect sizes are very small, showing that the distributions are nearly identical in practice. This supports our conclusion that the computational gains of g-DPO do not come at the expense of model quality.
Dataset Pair $D$ p-value Statistic location Sign
DhaA g-DPO - DPO 0.0290 <0.01<0.01 51.90 -1
DhaA g-DPO - Ref 0.1736 <0.01<0.01 52.49 -1
DhaA DPO - Ref 0.1851 <0.01<0.01 52.07 -1
Ty1 g-DPO - DPO 0.0385 <0.01<0.01 56.99 +1
Ty1 g-DPO - Ref 0.5325 <0.01<0.01 56.89 -1
Ty1 DPO - Ref 0.5695 <0.01<0.01 56.90 -1
Her2 g-DPO - DPO 0.0829 <0.01<0.01 0.2506 +1
Her2 g-DPO - Ref 0.3477 <0.01<0.01 0.2408 -1
Her2 DPO - Ref 0.3587 <0.01<0.01 0.2411 -1
Table 2: Two-sample Kolmogorov-Smirnov tests comparing predicted property distributions between models. The KS statistic $D$ reports the maximum distance between empirical CDFs; smaller values indicate closer distributions. $p$-values below $0.05$ indicate statistically significant differences.